Agenda/Minutes of the VIS Meeting on 2020/2/21
2020/2/21 14:00 - 16:00
Participants: Takahashi, Shoda, Tanaka, Akutsu, Fabian, Hirata, Aso
Progress report
Type-A
Past week report
- Enabled the one whitening filters for Oplevs for all Type-As.
 
Plan for coming weeks
Type B
Past week report
- The fiber scope was delivered. It had actually a head of 3.9mm in diameter. The fiber is a shape memory type.
 
Plan for coming weeks
Type-Bp
Past week report
- Enabled the two whitening filters for Oplevs for PR3 and PRM, and three for PR2.
 - Some roll offs have been set.
 
Plan for coming weeks
OMMT & OSTM
Past week report
Plan for coming weeks
VIS electronics
Past week report
- Measured the TFs from the output of ACT to the before/after demodulator.
 - The TFs should be compared with the theoretical calculation.
 
Plan for coming weeks
Other site works
Safety
Discussion
Uchiyama-san's schedule plan (linked by Shoda)
Travel Plans
- Travel (Week of 2/24): 
- Takahashi:
 - Sato:
 - Hirata:
 - Shoda:
 - Fujii:
 - Tanaka:
 
 - Travel (Week of 3/2): 
- Takahashi:
 - Sato:
 - Hirata:
 - Shoda:
 - Fujii:
 - Tanaka:
 
 
Next meeting
On 2020/2/28(Fri)
